### v1

@@ Abstract
Despite the recent developments in obstacle avoidance and other safety features, autonomous Unmanned Aerial Vehicles (UAVs) continue to face safety challenges. No previous work investigated the relationship between the behavioral uncertainty of a UAV and the unsafety of its flight. By quantifying uncertainty, it is possible to develop a predictor for unsafety, which acts as a flight supervisor. We conducted a large-scale empirical investigation of safety violations using PX4-Autopilot, an open-source UAV software platform. Our dataset of over 5,000 simulated flights, created to challenge obstacle avoidance, allowed us to explore the relation between uncertain UAV decisions and safety violations: up to 89% of unsafe UAV states exhibit significant decision uncertainty, and up to 74% of uncertain decisions lead to unsafe states. Based on these findings, we implemented Superialist (Supervising Autonomous Aerial Vehicles), a runtime uncertainty detector based on autoencoders, the state-of-the-art technology for anomaly detection. Superialist achieved high performance in detecting uncertain behaviors with up to 96% precision and 93% recall. Despite the observed performance degradation when using the same approach for predicting unsafety (up to 74% precision and 87% recall), Superialist enabled early prediction of unsafe states up to 50 seconds in advance.
@@ 1 Introduction
Autonomous Cyber-physical systems (CPSs), such as Unmanned Aerial Vehicles (UAVs), are employed in different application scenarios, including crop monitoring CarboneAMOSKNT21 and surveillance DBLP:journals/sensors/BalestrieriDVPT21 , to search and rescue in disaster areas ChowdhurySMLB21 , which pose significant challenges when ensuring UAV operational safety DAndrea14 ; ChowdhurySMLB21 ; UAV:2022 ; ZampettiKPP22 ; ICST-UAV2025 . UAV developers enhance the safety measures of the autonomous control modules in various ways, e.g., by incorporating functionalities for autonomous obstacle avoidance, collision prevention, safe takeoff, and landing, or by defining a Geo-fence (i.e., a virtual boundary of the safe flight zone). Despite such safety features being embedded in both commercial drones and popular open-access UAV Autopilot systems ardupilot ; px4-Autpilot , various incidents involving drones accidents:2019 suggest that UAVs present relevant safety challenges to be addressed by researcher and UAV developers.
Testing UAV functions and features in critical scenarios is still lacking. For this reason, researchers proposed to use simulation environments to test general CPSs and UAVs in a diversified set of scenarios HuangSLFB21 ; BojarczukGLDH0S21 ; PiazzoniCAYSV21 ; SDCScissor ; NguyenHG21 ; surrealist , and to support testing automation AlconTAC21 ; Wotawa21 , automated test generation surrealist , regression testing Birchler2022Machine ; abs-2107-09614 and debugging SmithR21 ; RoyHACC21 activities. However, while simulation-based testing remains a crucial practice for evaluating the functional and safety requirements before any real-world flight afzal2018crashing ; afzal2021simulation ; wang2021exploratory ; surrealist , even with the most advanced testing processes in place, UAVs may deviate from their intended behavior at runtime surrealist ; 3564821 ; 3571854 ; ZAMPETTI2022111425 due to many potentially unforeseen runtime factors such as weather conditions and sensor noise. Hence, the current commercial deployment of autonomous systems such as UAVs (legally) requires the presence of a human pilot/driver to actively monitor their safe operation and be prepared to assume control if necessary matternet .
@@
To answer this question, we conducted the first large-scale data-driven empirical investigation of UAV safety violations. We considered PX4-Avoidance px4-avoidance , an open-source and widely used autonomous obstacle avoidance system built on top of PX4-Autopilot px4-Autpilot , one of the most popular open-source UAV Autopilot systems and firmware, as a relevant use case. Employing evolutionary algorithms surrealist , we simulated UAV flights in diverse environments and flight missions, revealing potential misbehaviors, crashes, and safety violations. Our extensive dataset features about 5,000 simulated UAV flights, including unsafe and uncertain conditions. Manual analysis of a subset enabled exploration of the link between uncertain UAV decisions and safety violations.
@@ 2.2.1 Decision Uncertainty
Ensuring the safety of autonomous systems requires the identification of unfamiliar contexts by modeling their uncertainty mcallister2017concrete . We can generally distinguish between two sources of uncertainty that affect the system’s performance: Aleatoric uncertainty accounts for the uncertainty truly present in the nature of the operational domain and can be seen as the randomness and ambiguity in the surrounding environment or the incapability of completely sensing all the details of the environment kendall2017uncertainties . Epistemic Uncertainty, also known as model uncertainty, accounts for uncertainty in the ML model parameters, or more generally, in the control algorithm, which descends from lack of knowledge (e.g., inadequate training set). This type of uncertainty captures the system’s lack of knowledge about certain operational scenarios due to insufficient training (or implementation) kendall2017uncertainties . In our study, we only consider the epistemic uncertainty involved in the control algorithm of our UAV system and minimize the effects of aleatoric uncertainty by repeating the simulations multiple times and by measuring the average behavior.
Autonomous systems, including UAVs, are constantly making decisions at runtime with the movement control decisions being the most important and safety-critical ones. Hence, we define decision uncertainty as follows:
@@
Intuitively, we characterize the drone’s runtime decision uncertainty as unsteady and inconsistent path planning demonstrated in the frequent and abnormal changes in the control signals of the drone in a short period of time and specific segment of the flight. Figure 3 demonstrates the relation between the high-level UAV control signals (i.e., safe waypoints that the drone is instructed to follow; left picture) and the resulting flight trajectory (i.e., actual positions of the drone derived from sensors; right picture) during an autonomous mission ($A\rightarrow B\rightarrow F$) with the presence of two obstacles. While the control waypoints in the initial sections of the flight ($A\rightarrow B\rightarrow C$) are consistent and the resulting flight trajectory is closely following them, they get inconsistent and deviate from the optimal path as the drone enters the critical area around the obstacles ($C\rightarrow D\rightarrow E$). The resulting flight trajectory also diverges from the optimal path. While in this case the decision uncertainty (especially around $C$) is observable even in the resulting flight trajectory, it is generally much more evident and easily measured in the UAV control signals (i.e., safe waypoints).
@@ 3.2 Anomaly Detection in UAV Components
Our work complements previous studies by doing the first empirical study on the relationship between the uncertainty in UAV decisions and the safety aspects of the flight. Our anomaly detection method proposed for uncertainty detection differs from the above studies by being completely back-box and not using any internal variables (sensor reading, motor signals, GPS), relying only on high-level control commands from an external path-planing module. This makes our approach ideal for monitoring high-level path planning algorithms (e.g., obstacle avoidance) external to the main auto-pilot system. Additionally, our method is evaluated using faults observed by the obstacle avoidance module in specific obstacle allocations, not synthetic faults injected into sensor values.
@@ 3.3 Simulation-based Testing of Autonomous Systems
Previous studies have characterized the software bugs in autonomous systems including UAV Autopilot platforms GarciaF0AXC20 ; UAV:2022 ; ZAMPETTI2022111425 and investigated the challenges associated with testing these systems in realistic environments garcia2020simulation ; afzal2020study ; afzal2021automated ; surrealist ; WuXZL23 ; MaAY21 ; VR-study . Wang et al. afzal2020study ; afzal2021automated focused on testing challenges in robotic systems, emphasizing the complexity of test environments. Complementary to this, Lidval et al. lindvall2017metamorphic developed an automated testing framework for drones in simulation to address the test oracle problem. Woodlief et al.introduced PHYS-FUZZ elbaum2021fuzzing , a fuzzing approach for mobile robots that considers their physical attributes and hazards. Hildebrandt et al. elbaum2021world proposed a "world-in-the-loop" mixed-reality approach for UAV testing that integrates sensor data from both simulated and real-world sources. In our previous work surrealist , we addressed the problem of realism in simulation-based testing for UAVs by proposing Surrealist, an approach that faithfully replicates real-world test scenarios in simulation environments and generates more challenging test case variants. Complementary, we developed a UAV test bench and test generation platform called Aerialist  aerialist , and on top of it we fostered research in this domain by organizing the first UAV test generation competition at the Intl. Workshop on Search-Based and Fuzz Testing (SBFT) uav-competition .
@@ 5.2 RQ2 - Uncertainty Detection
Since there is no established standard or guideline on the accepted or normal level of decision uncertainty for UAVs, we need to empirically determine a threshold based on our datasets. Hence, we rely on the histogram of the reconstruction loss derived from the nominal training data. Figure 7 visually represents the frequency distribution of different loss levels observed during nominal UAV operations. For the majority of time windows, the reconstruction loss is below 0.05, as depicted in the figure (note the logarithmic scale). However, there is a significant drop in the count for values exceeding 0.2, with the maximum value reaching 0.4. Given the expectation of a small number of anomalies in the training data that our filter may not have identified, we have decided to set the anomaly detection (i.e., uncertainty detection) threshold at $\theta=0.3$. Moreover, given the importance and criticality of early detection of uncertain decisions for preventing unsafe and crashing scenarios, we are mostly interested in increasing the True Positive Rate (i.e., correctly detected uncertainties) of the approach while accepting the associated possibility of increasing its False Positive Rate (i.e., false alarms).
@@ 5.4 Implications and Lessons Learned
The relationship between uncertainty and safety in UAVs is multifaceted. Safety, influenced by various factors including proximity to the surrounding obstacles and environmental conditions, does not always correlate with UAV uncertainty states. Thus, our proposed uncertainty estimation approach may not always directly address safety in UAV test scenarios. Nevertheless, an uncertainty-based monitoring system with a wide reaction window can still play a crucial role in enhancing the overall system safety by triggering autonomous self-healing mechanisms to recover the drone or alerting human pilots to potential risks, even if intervention is not always necessary. This is even more crucial considering the current regulations in commercial deployment of autonomous UAVs (e.g., for delivery tasks) that mandates the manual monitoring of the flights by remote pilots, where each pilot may monitor up to 20 concurrent flights matternet . Figure 9 demonstrates our suggested deployment architecture to use Superialist (after training the anomaly detection model) on board of PX4-powered UAVs, enabling sending real-time safety warnings to the autopilot system and the remote pilots.
@@ 5.5.2 External Validity
The datasets used for training and testing our model were derived from simulated UAV flights. The extent to which our findings can be generalized to real-world UAV operations is uncertain. The simulation may not fully capture variations in environmental conditions, UAV models, and operational scenarios in the real world. However, simulation-based testing is considered a best practice before moving to field testing. Moreover, the range of flight scenarios and obstacle configurations in our simulations may not represent all possible real-world challenges that UAVs might encounter. This limitation could affect the applicability of our results to more complex or unforeseen flight conditions.
@@ 6 Conclusion and Future Work
In this study, we explored the relationship between decision uncertainty and flight safety in UAV operations. Our empirical study focused on real-time detection of uncertain and unsafe behaviors, through the analysis of simulated UAV flight data. Key findings indicate a moderate correlation between uncertain decision-making and unsafe flight outcomes, with the Auto-encoder model demonstrating high precision and recall in uncertainty detection. Despite the lower precision in unsafety prediction, our safety monitoring approach with a wide reaction window can help to enhance safety in UAV operations by triggering fail-safe mechanisms or alerting human pilots.
This research not only sheds light on the critical aspects of UAV safety but also opens ways for more advanced, data-driven safety measures in autonomous flight systems. In future work, we aim to focus on identifying additional features, such as white/gray box indicators of uncertainty, to enhance unsafety prediction. For instance, white box indicators of decision uncertainty could be obtained by inserting probes into the obstacle avoidance module. Moreover, we target to extend and replicate our study in field testing scenarios as well as other (e.g., robotics) domains.
Acknowledgements.
We thank the Horizon 2020 (EU Commission) support for the project InnoGuard, Marie Sktodowska-Curie Actions Doctoral Networks (HORIZON-MSCA-2023-DN), and the Hasler Foundation for the project (Project No.23064) entitled "Bridging the Reality Gap in Testing Unmanned Aerial Vehicles".
Data Availability
@@ Acknowledgements.
Declarations
Funding and/or Conflicts of interests/Competing interests.

### v2

@@ Abstract
Despite the recent developments in obstacle avoidance and other safety features, autonomous Unmanned Aerial Vehicles (UAVs) continue to face safety challenges. No previous work investigated the relationship between the behavioral uncertainty of a UAV, characterized in this work by inconsistent or erratic control signal patterns, and the unsafety of its flight. By quantifying uncertainty, it is possible to develop a predictor for unsafety, which acts as a flight supervisor. We conducted a large-scale empirical investigation of safety violations using PX4-Autopilot, an open-source UAV software platform. Our dataset of over 5,000 simulated flights, created to challenge obstacle avoidance, allowed us to explore the relation between uncertain UAV decisions and safety violations: up to 89% of unsafe UAV states exhibit significant decision uncertainty, and up to 74% of uncertain decisions lead to unsafe states. Based on these findings, we implemented Superialist (Supervising Autonomous Aerial Vehicles), a runtime uncertainty detector based on autoencoders, the state-of-the-art technology for anomaly detection. Superialist achieved high performance in detecting uncertain behaviors with up to 96% precision and 93% recall. Despite the observed performance degradation when using the same approach for predicting unsafety (up to 74% precision and 87% recall), Superialist enabled early prediction of unsafe states up to 50 seconds in advance.
@@ 1 Introduction
Autonomous Cyber-physical systems (CPSs), such as Unmanned Aerial Vehicles (UAVs), are employed in different application scenarios, including crop monitoring CarboneAMOSKNT21 and surveillance DBLP:journals/sensors/BalestrieriDVPT21 , to search and rescue in disaster areas ChowdhurySMLB21 , which pose significant challenges when ensuring UAV operational safety DAndrea14 ; ChowdhurySMLB21 ; UAV:2022 ; ZAMPETTI2022111425 ; ICST-UAV2025 . UAV developers enhance the safety measures of the autonomous control modules in various ways, e.g., by incorporating functionalities for autonomous obstacle avoidance, collision prevention, safe takeoff, and landing, or by defining a Geo-fence (i.e., a virtual boundary of the safe flight zone). Despite such safety features being embedded in both commercial drones and popular open-access UAV Autopilot systems ardupilot ; px4-Autpilot , various incidents involving drones accidents:2019 suggest that UAVs present relevant safety challenges to be addressed by researcher and UAV developers 3711906 .
Testing UAV functions and features in critical scenarios is still lacking. For this reason, researchers proposed to use simulation environments to test general CPSs and UAVs in a diversified set of scenarios HuangSLFB21 ; BojarczukGLDH0S21 ; PiazzoniCAYSV21 ; SDCScissor ; NguyenHG21 ; surrealist , and to support testing automation AlconTAC21 ; Wotawa21 , automated test generation surrealist , regression testing Birchler2022Machine ; Birchler2022Single and debugging SmithR21 ; RoyHACC21 activities. However, while simulation-based testing remains a crucial practice for evaluating the functional and safety requirements before any real-world flight afzal2018crashing ; afzal2021simulation ; wang2021exploratory ; surrealist , even with the most advanced testing processes in place, UAVs may deviate from their intended behavior at runtime surrealist ; 3571854 ; ZAMPETTI2022111425 ; UAV:2022 due to many potentially unforeseen runtime factors such as weather conditions and sensor noise. Hence, the current commercial deployment of autonomous systems such as UAVs (legally) requires the presence of a human pilot/driver to actively monitor their safe operation and be prepared to assume control if necessary matternet .
@@
To answer this question, we conducted the first large-scale data-driven empirical investigation of UAV safety violations and their relation to decision uncertainty. We considered PX4-Avoidance px4-avoidance , an open-source and widely used autonomous obstacle avoidance system built on top of PX4-Autopilot px4-Autpilot , one of the most popular open-source UAV Autopilot systems and firmware, as a relevant use case. Employing evolutionary algorithms surrealist , we simulated UAV flights in diverse environments and flight missions, revealing potential misbehaviors, crashes, and safety violations. Our extensive dataset features about 5,000 simulated UAV flights, including unsafe and uncertain conditions. Manual analysis of a subset enabled exploration of the link between uncertain UAV decisions and safety violations.
@@ 2.2.1 Decision Uncertainty
Ensuring the safety of autonomous systems requires the identification of unfamiliar contexts by modeling their uncertainty mcallister2017concrete . We can generally distinguish between two sources of uncertainty that affect the system’s performance: Aleatoric uncertainty accounts for the uncertainty truly present in the nature of the operational domain and can be seen as the randomness and ambiguity in the surrounding environment or the incapability of completely sensing all the details of the environment kendall2017uncertainties . Epistemic Uncertainty, also known as model uncertainty, accounts for uncertainty in the ML model parameters, or more generally, in the control algorithm, which descends from lack of knowledge (e.g., inadequate training set). This type of uncertainty captures the system’s lack of knowledge about certain operational scenarios due to insufficient training (or implementation) kendall2017uncertainties . In safety engineering —the application of engineering principles and methods to identify, analyze, and control hazards within a system throughout its lifecycle to achieve an acceptable level of safety and prevent harm— uncertainty often refers to the lack of knowledge about factors contributing to risk (i.e., epistemic uncertainty), such as the precise failure rate of components, the exact environmental conditions, or the completeness of system models ryan2024dynamic ; denney2015dynamic ; hawkins2013assurance .
While we acknowledge this established view, our study focuses on a different, albeit related, aspect: the observable behaviors arising when an autonomous system encounters situations challenging its control logic, potentially stemming from underlying epistemic or aleatoric uncertainty. We name this phenomenon Behavioral Control Uncertainty (or simply Decision Uncertainty within the context of our study – to clarify the differences and commonalities between this definition and the existing ones) and define it operationally based on the system’s output:
@@
Intuitively, we characterize this behavioral uncertainty by analyzing the stability and consistency of the drone’s planned path (control signals), rather than directly measuring uncertainty about internal parameters or external risk factors. Our rationale for this black-box, output-focused approach is its potential for real-time monitoring, using readily available control data, and its ability to capture system issues in complex scenarios, regardless of the specific internal cause. This study empirically investigates the extent to which this observable behavioral uncertainty correlates with, and potentially predicts, unsafe flight states.
Figure 3 demonstrates the relation between the high-level UAV control signals (i.e., safe waypoints that the drone is instructed to follow; left picture) and the resulting flight trajectory (i.e., actual positions of the drone derived from sensors; right picture) during an autonomous mission ($A\rightarrow B\rightarrow F$) with the presence of two obstacles. While the control waypoints in the initial sections of the flight ($A\rightarrow B\rightarrow C$) are consistent and the resulting flight trajectory is closely following them, they get inconsistent and deviate from the optimal path as the drone enters the critical area around the obstacles ($C\rightarrow D\rightarrow E$). The resulting flight trajectory also diverges from the optimal path. While in this case the decision uncertainty (especially around $C$) is observable even in the resulting flight trajectory, it is generally much more evident and easily measured in the UAV control signals (i.e., safe waypoints).
@@ 3 Related Works
3.1 Safety Assurance for UAVs
The traditional aviation industry adheres to well-established and rigorous safety standards governing system development and safety assessment processes SAE:ARP4754A ; SAE:ARP4761 . These standards mandate comprehensive design assurance and verification activities scaled according to potential risks, forming the basis for demonstrating aircraft safety. In contrast, the rapidly evolving field of consumer and small UAVs currently operates under less comprehensive and standardized design assurance frameworks. While regulations —such as the EASA drone classification system (C0-C4) EASA:Reg2019_945 , drone operational categories (Open, Specific, Certified) EASA:Reg2019_947 , and CE marking requirements EU:RED2014_53 )— exist concerning specific technical capabilities, operational limitations, and general product safety for consumer drones EUDronePort:C1Reqs , they typically do not enforce the same depth of system development and safety assessment rigor found in traditional aviation. This difference highlights a potential gap and an ongoing challenge in establishing and demonstrating sufficient safety assurance for autonomous UAVs, particularly as their operational complexity increases and they interact more with dynamic environments.
Safety assurance is the process of providing justified confidence that a system meets its defined safety requirements and is acceptably safe for its intended purpose within a specific operational context hawkins2013assurance . Traditionally, this process leads to the creation of a safety case: a structured argument, supported by a body of evidence (e.g., analysis results, test reports, simulations), demonstrating that the system’s risks are tolerable hawkins2013assurance . Safety cases typically reflect the system understanding and safety assessment performed at design time. However, for autonomous systems such as UAVs operating in dynamic and unpredictable environments, this static design-time assurance may become insufficient or invalidated during operation as underlying assumptions change ryan2024dynamic ; denney2015dynamic .
To address this challenge, the concept of Dynamic Safety Assurance (DSA) has emerged, advocating for continuous, through-life safety management  ryan2024dynamic ; denney2015dynamic . DSA frameworks emphasize the need to monitor the system and its operational context at runtime, continuously assess risks and the validity of safety arguments, and potentially adapt the system behavior to maintain safety. Mechanisms like Dynamic Safety Cases (DSCs) have been proposed to structure this process, integrating runtime evidence to dynamically update the estimated confidence levels in safety claims ryan2024dynamic . A cornerstone of DSA is the development of effective runtime monitors capable of detecting anomalies, performance degradation, or emerging hazards ryan2024dynamic . Our work contributes to this area by investigating decision uncertainty as a runtime indicator of potential safety issues and proposing SUPERIALIST, a black-box monitor designed to detect such uncertainty, thereby providing evidence that could be leveraged within a DSA framework.
@@ 3.3 Anomaly Detection in UAV Components
Our work complements previous studies by doing the first empirical study on the relationship between the uncertainty in UAV decisions and the safety aspects of the flight. Our anomaly detection method proposed for uncertainty detection differs from the above studies by being completely back-box and not using any internal variables (sensor reading, motor signals, GPS), relying only on high-level control commands from an external path planing module. This makes our approach ideal for monitoring high-level path planning algorithms (e.g., obstacle avoidance) external to the main auto-pilot system. Additionally, our method is evaluated using faults observed by the obstacle avoidance module in specific obstacle allocations, not synthetic faults injected into sensor values.
@@ 3.4 Simulation-based Testing of Autonomous Systems
Previous studies have characterized the software bugs in autonomous systems including UAV Autopilot platforms GarciaF0AXC20 ; UAV:2022 ; ZAMPETTI2022111425 and investigated the challenges associated with testing these systems in realistic environments garcia2020simulation ; afzal2020study ; surrealist ; WuXZL23 ; MaAY21 ; VR-study . Afzal et al. afzal2020study focused on testing challenges in robotic systems, emphasizing the complexity of test environments. Complementary to this, Lidval et al. lindvall2017metamorphic and Afzal et al. afzal2021mithra developed automated testing frameworks for drones in simulation to address the test oracle problem. Woodlief et al. introduced PHYS-FUZZ elbaum2021fuzzing , a fuzzing approach for mobile robots that considers their physical attributes and hazards. Hildebrandt et al. elbaum2021world proposed a "world-in-the-loop" mixed-reality approach for UAV testing that integrates sensor data from both simulated and real-world sources. In our previous work surrealist , we addressed the problem of realism in simulation-based testing for UAVs by proposing Surrealist, an approach that faithfully replicates real-world test scenarios in simulation environments and generates more challenging test case variants. Complementary, we developed a UAV test bench and test generation platform called Aerialist  aerialist , and on top of it we fostered research in this domain by organizing the first UAV test generation competition at the Intl. Workshop on Search-Based and Fuzz Testing (SBFT) uav-competition ; ICST-UAV2025 ; SBFT-UAV2025 .
@@ 5.2 RQ2 - Uncertainty Detection
Since there is no established standard or guideline on the accepted or normal level of decision uncertainty for UAVs, we need to empirically determine a threshold based on our datasets. Hence, we rely on the histogram of the reconstruction loss derived from the nominal training data. Figure 7 visually represents the frequency distribution of different loss levels observed during nominal UAV operations. For the majority of time windows, the reconstruction loss is below 0.05, as depicted in the figure (note the logarithmic scale). However, there is a significant drop in the count for values exceeding 0.2, with the maximum value reaching 0.4. Given the expectation of a small number of anomalies in the training data that our filter may not have identified, we have decided to set the anomaly detection (i.e., uncertainty detection) threshold at $\theta=0.3$. Moreover, given the importance and criticality of early detection of uncertain decisions for preventing unsafe and crashing scenarios, we are interested in increasing the True Positive Rate (i.e., correctly detected uncertainties) of the approach while accepting a limited possibility of increasing its False Positive Rate (i.e., false alarms) as a consequence.
@@ 5.4 Implications and Lessons Learned
The relationship between uncertainty and safety in UAVs is multifaceted. Safety, influenced by various factors including proximity to the surrounding obstacles and environmental conditions, does not always correlate with UAV uncertainty states. Thus, our proposed uncertainty estimation approach may not always directly address safety in UAV test scenarios. Nevertheless, an uncertainty-based monitoring system with a wide reaction window can still play a crucial role in enhancing the overall system safety by triggering autonomous self-healing mechanisms to recover the drone or alerting human pilots to potential risks, even if intervention is not always necessary. This is even more crucial considering the current regulations in commercial deployment of autonomous UAVs (e.g., for delivery tasks) that mandates the manual monitoring of the flights by remote pilots, where each pilot may monitor up to 20 concurrent flights matternet . Figure 9 demonstrates our suggested deployment architecture to use Superialist (after training the anomaly detection model) on board of PX4-powered UAVs, enabling sending real-time safety warnings to the autopilot system and the remote pilots. This signal can also be integrated in a dynamic safety assurance framework ryan2024dynamic and trigger configuration changes to maintain safety of the system.
@@ 5.5.2 External Validity
The datasets used for training and testing our model were derived from simulated UAV flights. The extent to which our findings can be generalized to real-world UAV operations is uncertain. The simulation may not fully capture variations in environmental conditions, UAV models, and operational scenarios in the real world. However, simulation-based testing is considered a best practice before moving to field testing. Moreover, the range of flight scenarios, environmental conditions, and obstacle configurations in our simulations may not represent all possible real-world challenges that UAVs might encounter. More specifically, for simplicity, we only considered a flight mission with three waypoints, two box-shaped obstacles, and without any wind. This limitation could affect the applicability of our results to more complex or unforeseen flight conditions, and to uncertain control decisions due to factors other than obstacles (e.g., wind).
@@ 6 Conclusion and Future Work
In this study, we explored the relationship between decision uncertainty and flight safety in UAV operations. Our empirical study focused on real-time detection of uncertain and unsafe behaviors, through the analysis of simulated UAV flight data. Key findings indicate a moderate correlation between uncertain decision-making and unsafe flight outcomes, with the Auto-encoder model demonstrating high precision and recall in uncertainty detection. Despite the lower precision in unsafety prediction, our safety monitoring approach with a wide reaction window can help to enhance safety in UAV operations by triggering fail-safe mechanisms, alerting human pilots, or integration into a dynamic safety assurance framework.
This research not only sheds light on the critical aspects of UAV safety but also opens ways for more advanced, data-driven safety measures in autonomous flight systems. In future work, we aim to focus on identifying additional features, such as drone’s speed and white/gray box indicators of uncertainty, to enhance unsafety prediction. For instance, white box indicators of decision uncertainty could be obtained by inserting probes into the obstacle avoidance module. Moreover, we aim to extend and further evaluate our approach in dynamic environments (i.e., including moving obstacles), windy weather, in the field (i.e., real-world tests), and in other robotic domains.
Declarations
Funding:
We thank the Horizon 2020 (EU Commission) support for the project InnoGuard, Marie Sktodowska-Curie Actions Doctoral Networks (HORIZON-MSCA-2023-DN), and the SNSF project entitled "SwarmOps: Human-sensing based MLOps for Collaborative Cyber-physical systems" (Project No. 200021_219732). We also thank the Hasler Foundation for the projects "Bridging the Reality Gap in Testing Unmanned Aerial Vehicles" (Project No.23064) and "Advancing Unmanned Aerial Vehicles Reliability and Societal Trust Through Integrated Testing and Formal Verification" (Project No.2025-02-27-311).
Data Availability:
@@
Author Contributions:
  • Sajad Khatiri: Conceptualization, Methodology, Software, Data Curation, Validation, Investigation, Writing – Original Draft
  • Fatemeh Mohammadi Amin: Methodology, Software, Validation, Data Curation, Visualization, Writing – Original Draft
  • Sebastiano Panichella: Conceptualization, Methodology, Data Curation, Writing – Review & Editing, Supervision, Project Administration, Funding Acquisition
  • Paolo Tonella: Conceptualization, Methodology, Writing – Review & Editing, Supervision
Conflict of Interest:
@@
Ethical Approval:
not applicable.
Informed Consent:
not applicable.
Clinical Trial Number:
not applicable.